what two consecutive even integers whose sum is 400?
Answer by 24HoursTutor.com(40)

Consecutive numbers mean numbers which come right after one another.For ex 3 and 4, 45 and 46, etc.
Even numbers are number which are divisible by 2.
So, consecutive even numbers mean numbers which come after one another and are divisible by 2 for eg. 4 and 6, 22 and 24, etc.
Let us assume the first number to be x. For the other number to be consecutive it has to be 2 more than x. So, our second number is x + 2. Now we are given the sum of these two numbers is 400. So, our equation would look something like this :
x+x+2 = 400
2x + 2=400
2x= 400-2
2x= 398
x=398/2
x=199
and second number is x+2 = 199+2 = 201
The two consecutive numbers which form a sum of 400 are 199 and 201 but they are odd. There are no two even consecutive numbers which can form a sum of 400.
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr